OpenStudy (anonymous):

Judith can ride her bike 2 mi/h faster than kelly. Judith takes 48 min less to travel 50 miles than Kelly does. What is each person's speed in miles per hour??

OpenStudy (dumbcow):

Let j be judiths speed and k be kellys speed j = k+2 we also know distance = rate*time convert 48 min to hours (48/60 = 4/5) judith: 50 = (t-4/5)*j kelly: 50 = t*k set both equations equal to each other and replace j with (k+2) tk = (t-4/5)(k+2) tk = tk +2t -4/5k -8/5 0=2t -4/5k -8/5 t = 2/5k + 4/5 substitute this value in for t to find k 50 = k(2/5k +4/5) 2/5k^2 +4/5k -50 = 0 get rid of fractions by multiplying equation by 5 2k^2 +4k -250 =0 quadratic equation gives k = 10.225 mph j = k +2 j = 12.225 mph